# Weak compactness cardinals for strong logics and subtlety properties of the class of ordinals

### v1

@@ 5. Abstract logics
      If $f\in{\mathfrak{F}}_{\tau}$ and $x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(f)}\in{{\rm{dom}}(M)}$, then
      $\pi(f^{M}(x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(f)}))~{}=~{}f^{N}(\pi(x_{0}),\ldots,\pi(x_{{\mathfrak{a}}_{\tau}(f)})).$
@@
      If $r\in{\mathfrak{R}}_{\tau}$ and $x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(r)}\in{{\rm{dom}}(M)}$, then
      $r^{M}(x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(r)})~{}\Longleftrightarrow~{}r^{N}(\pi(x_{0}),\ldots,\pi(x_{{\mathfrak{a}}_{\tau}(r)})).$
@@
    A language $\sigma$ is a sublanguage of a language $\tau$ if $\mathfrak{C}_{\sigma}\subseteq\mathfrak{C}_{\tau}$, $\mathfrak{F}_{\sigma}\subseteq\mathfrak{F}_{\tau}$, $\mathfrak{R}_{\sigma}\subseteq\mathfrak{R}_{\tau}$ and $a_{\sigma}=a_{\tau}\restriction(\mathfrak{F}_{\sigma}\cup\mathfrak{R}_{\sigma})$.
@@
    satisfying $u[\mathfrak{C}_{\sigma}]=\mathfrak{C}_{\tau}$, $u[\mathfrak{F}_{\sigma}]=\mathfrak{F}_{\tau}$, $u[\mathfrak{R}_{\sigma}]=\mathfrak{R}_{\tau}$ and $a_{\sigma}(s)=a_{\tau}(u(s))$ for all $s\in\mathfrak{F}_{\sigma}\cup\mathfrak{R}_{\sigma}$.
@@
In [5, Section 4], it is verified that Makowsky’s Theorem 1.1 holds true with respect to the above definitions. In another direction, an unpublished result of Stavi shows that the validity of Vopěnka’s Principle is also equivalent to the fact that analogs of the Löwenheim–Skolem Downward Theorem holds for all abstract logics (see [15, Theorem 6]). For sake of completeness and to motivate the notions introduced in items (8) and (9) of Definition 5.3, we present a proof of Stavi’s result.
@@
  • Given is a language $\tau$, $M\in Str_{\tau}$ and a binary relation symbol $r$ in $\tau$, then $M\models_{\mathcal{L}_{\Gamma}}\phi_{r}$ holds if and only if $\langle\bar{M},r^{M}\rangle$ is a graph that is isomorphic to a graph in $\Gamma$.
Our assumption now yields an uncountable cardinal $\kappa$ with the property that for every language $\tau$ with $|{{\mathfrak{S}}_{\tau}}|<\kappa$, every $\tau$-structure $N$ and every $\phi\in\mathcal{L}_{\Gamma}(\tau)$ with $N\models_{\mathcal{L}_{\Gamma}}\phi$, there exists a substructure $M$ of $N$ with $|{{{\rm{dom}}(M)}}|<\kappa$ and $M\models_{\mathcal{L}_{\Gamma}}\phi$. Since $\Gamma$ is a proper class, we can fix a graph $G$ in $\Gamma$ that has cardinality greater than $\kappa$. Let $\sigma$ denote a language with a single binary relation symbol $e$ and let $\tau$ be a language extending $\sigma$ that adds functions symbols for first-order Skolem functions to $\sigma$. View $G$ as a $\sigma$-structure and let $G^{\prime}$ be a $\tau$-structure whose $\sigma$-reduct is $G$ and that interprets the new function symbols as the corresponding first-order Skolem functions. Then item (4) in Definition 5.2 ensures that $G^{\prime}\models_{\mathcal{L}_{\Gamma}}\phi_{e}$ holds and, since $|{{\mathfrak{S}}_{\tau}}|<\kappa$, we can use our assumption to find a substructure substructure $H^{\prime}$ of $G^{\prime}$ with $|{H^{\prime}}|<\kappa$ and $H^{\prime}\models_{\mathcal{L}_{\Gamma}}\phi_{e}$. Let $H$ denote the $\sigma$-reduct of $H^{\prime}$. Our setup now ensures that $H$ is a first-order elementary submodel of $G$ and $H\models_{\mathcal{L}_{\Gamma}}\phi_{e}$ holds. This shows that $\langle{{\rm{dom}}(H)},e^{H}\rangle$ is a graph that is isomorphic to an element $K$ of $\Gamma$. Since ${{\rm{dom}}(H)}$ has cardinality less than $\kappa$, we know that $K\neq G$ and $K$ embeds into $G$. These computations show that (1) holds.
@@ 6. Structural properties of abstract logics
(1) Note that, given a language $\sigma$, an application of (6) in Definition 5.2 to the trivial renaming of $\sigma$ into itself ensures that the identity function on $\mathcal{L}(\sigma)$ is the unique permutation $\pi$ on $\mathcal{L}(\sigma)$ with the property that
$M\models_{\mathcal{L}}\phi~{}\Longleftrightarrow~{}M\models_{\mathcal{L}}\pi(\phi)$
@@
(2) Let $\mu$ be a cardinal as in (1), let $\sigma$ be a language in ${\rm{H}}(o_{\mathcal{L}})$, let $u$ be a renaming of $\sigma$ into a language $\tau$ and let $\varpi$ be a bijection $\pi$ between $\mathcal{L}(\sigma)$ and $\mathcal{L}(\tau)$ such that (C) holds. Assume, towards a contradiction that $\varpi\neq u_{*}$. Then $\pi=(u_{*})^{{-}1}\circ\varpi$ is a non-trivial permutation of $\mathcal{L}(\sigma)$ and, since $M_{\sigma,\pi}$ is an element of ${\rm{H}}(\mu)$, our assumptions on $\varpi$ ensure that
@@
holds for all $M\in Str_{\tau_{1}}\cap X$ and all $\phi\in\mathcal{L}(\tau_{1})$. Since ${\rm{H}}(\mu)\subseteq X$, we can now apply (2) to conclude that $\varpi=u_{*}$
@@
(4) Let $\mu$ be a cardinal as in (1), let $\kappa>\mu$ be a $C^{(n)}$-weakly shrewd cardinal and let $U$ be a ${<}\kappa$-satisfiable theory of cardinality $\kappa$. Since $o_{\mathcal{L}}<\kappa$, we can find a language $\sigma$ with $|{{\mathfrak{S}}_{\sigma}}|\leq\kappa$ and $U\subseteq\mathcal{L}(\sigma)$. Then there is a renaming $u$ of $\sigma$ into a language $\tau$ with ${\mathfrak{S}}_{\tau}\subseteq{\rm{H}}(\kappa)$. Set $T=u_{*}[U]$. Then (6) in Definition 5.2 ensures that $T$ is also ${<}\kappa$-satisfiable. Now, pick a cardinal $\kappa<\theta\in C^{(n+2)}$ with $T\in{\rm{H}}(\theta)$ and use (3) to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\mu<\bar{\kappa}<\min(\kappa,\bar{\theta})$, a sublanguage $\bar{\tau}$ of $\tau$ with $|{{\mathfrak{S}}_{\bar{\tau}}}|\leq\bar{\kappa}$ and ${\mathfrak{S}}_{\bar{\tau}}={\mathfrak{S}}_{\tau}\cap{\rm{H}}(\bar{\kappa})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}\cup\{\bar{\kappa},\bar{\tau},\mathcal{L}(\bar{\tau}),z\}\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$, $j(z)=z$, $j(\bar{\tau})=\tau$, $j(\mathcal{L}(\bar{\tau}))=\mathcal{L}(\tau)$, $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$ and $T\in{{\rm{ran}}(j)}$. Pick $\bar{T}\in X$ with $j(\bar{T})=T$. Then $\bar{T}$ is a subset of $\mathcal{L}(\bar{\tau})$ of cardinality $\bar{\kappa}$ and the fact that $\bar{\kappa}\subseteq X$ ensures that $\bar{T}\subseteq X$. Moreover, since $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$, it follows that $\bar{T}=j[\bar{T}]\subseteq T$ and hence $\bar{T}$ is satisfiable. This shows that there is a $\bar{\tau}$-structure $M$ with the property that $\psi(M,\phi,z)$ holds for all $\phi$ in $\bar{T}$. The correctness properties of ${\rm{H}}(\bar{\theta})$ now ensure that this statement holds in $X$ and hence we know that, in ${\rm{H}}(\theta)$, there is a $\tau$-structure $N$ with the property that $\psi(N,\phi,z)$ holds for all $\phi$ in $T$. Using the correctness properties of ${\rm{H}}(\theta)$, we can now conclude that $T$ is satisfiable and, by (6) in Definition 5.2, we also know that $U$ is satisfiable.
@@
(5) Let $\mu$ be a cardinal as in (1), let $\kappa>\mu$ be a $C^{(n)}$-weakly shrewd cardinal, let $\sigma$ be a language with the property that $|{{\mathfrak{S}}_{\sigma}}|<\kappa$, let $I$ be a $\sigma$-structure with $|{{{\rm{dom}}(I)}}|=\kappa$ and let $\phi$ be an element of $\mathcal{L}(\sigma)$ with $I\models_{\mathcal{L}}\phi$. Let $u$ be a renaming of $\sigma$ into a language $\tau\in{\rm{H}}(\kappa)$. Set $K=u^{*}(I)$ and $\phi_{1}=u_{*}(\phi)$. Then $K\models_{\mathcal{L}}\phi_{1}$ and there is an isomorphism $\pi$ between $K$ and a $\tau$-structure $N$ with ${{\rm{dom}}(N)}=\kappa$ and $N\models_{\mathcal{L}}\phi_{1}$. Fix a cardinal $\kappa<\theta\in C^{(n+2)}$. Then $K$ is an element of ${\rm{H}}(\theta)$ and we can use (3) to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\mu<\bar{\kappa}<\min(\kappa,\bar{\theta})$, a sublanguage $\bar{\tau}$ of $\tau$ with $|{{\mathfrak{S}}_{\bar{\tau}}}|\leq\bar{\kappa}$ and ${\mathfrak{S}}_{\bar{\tau}}={\mathfrak{S}}_{\tau}\cap{\rm{H}}(\bar{\kappa})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}\cup\{\bar{\kappa},\bar{\tau},\mathcal{L}(\bar{\tau}),z\}\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$, $j(z)=z$, $j(\bar{\tau})=\tau$, $j(\mathcal{L}(\bar{\tau}))=\mathcal{L}(\tau)$, $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$ and $N,\phi_{1}\in{{\rm{ran}}(j)}$. The fact that $\tau$ is an element of ${\rm{H}}(\kappa)$ then implies that $\tau=\bar{\tau}\in{\rm{H}}(\bar{\kappa})$, ${\mathfrak{S}}_{\tau}\subseteq X$ and $j\restriction{\mathfrak{S}}_{\tau}={\rm{id}}_{{\mathfrak{S}}_{\tau}}$. Pick $M$ in $X$ with $j(M)=N$. Then $M$ is a $\tau$-structure with ${{\rm{dom}}(M)}=\bar{\kappa}$ and it follows that $M$ is a substructure of $N$. Next, pick $\phi_{0}$ in $X$ with $j(\phi_{0})=\phi_{1}$. Then $\phi_{0}\in\mathcal{L}(\tau)$ and the fact that $j\restriction(\mathcal{L}(\tau)\cap X)={\rm{id}}_{\mathcal{L}(\tau)\cap X}$ implies that $\phi_{0}=\phi_{1}\in X$. Moreover, using the elementarity of $j$ and the correctness properties of ${\rm{H}}(\theta)$ and ${\rm{H}}(\bar{\theta})$, we know that $M\models_{\mathcal{L}}\phi_{1}$ holds. Then $\pi^{{-}1}\restriction\bar{\kappa}$ is an isomorphism between $M$ and a substructure $J$ of $K$, and we can apply (5) in Definition 5.2 to show that $J\models_{\mathcal{L}}\phi_{1}$. Finally, we can use (5) in Definition 5.2 to conclude that $H=(u^{*})^{{-}1}(J)$ is a substructure of $I$ with $|{{{\rm{dom}}(H)}}|<\kappa$ and $H\models_{\mathcal{L}}\phi$.
@@
Fix $\phi\in\mathcal{L}(\tau)$. By (7) in Definition 5.2, there is a sublanguage $\tau_{0}$ of $\tau$ with $|{{\mathfrak{S}}_{\tau_{0}}}|<o_{\mathcal{L}}$ and $\phi\in\mathcal{L}(\tau_{0})$. Since $|{{\mathfrak{S}}_{\tau}}|^{{<}o_{\mathcal{L}}}\leq\bar{\kappa}\subseteq X$, it follows that $\tau_{0}$ is an element of ${\rm{H}}(\bar{\kappa})\cap X$ and this implies that $j(\tau_{0})=\tau_{0}$ and $j(\mathcal{L}(\tau_{0}))=\mathcal{L}(\tau_{0})$. Now, we can find a language $\tau_{1}\in{\rm{H}}(o_{\mathcal{L}})$ and a renaming $v$ of $\tau_{1}$ into $\tau_{0}$ that is an element of $X$. Our setup then ensures that $\mathcal{L}(\tau_{1})\in{\rm{H}}(\mu)\subseteq X$ and $j(\mathcal{L}(\tau_{1}))=\mathcal{L}(\tau_{1})$. Since (2) shows that $v_{*}$ is the unique bijection between $\mathcal{L}(\tau_{1})$ and $\mathcal{L}(\tau_{0})$ with the property that (C) holds for all $\tau_{1}$-structures $M$ in ${\rm{H}}(\mu)$ and every $\phi\in\mathcal{L}(\tau_{1})$, the fact that all parameters of this existence statements are elements of $X$ that are fixed by $j$ ensures that $v_{*}$ is an element of $X$. In particular, since $\mathcal{L}(\tau_{1})\subseteq{\rm{H}}(\mu)\subseteq X$, we know that $\mathcal{L}(\tau_{0})\subseteq X$ and hence $\phi$ is an element of $X$. ∎

### v3

@@ 5. Abstract logics
      If $f\in{\mathfrak{F}}_{\tau}$ and $x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(f)-1}\in{{\rm{dom}}(M)}$, then
      $\pi(f^{M}(x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(f)-1}))~{}=~{}f^{N}(\pi(x_{0}),\ldots,\pi(x_{{\mathfrak{a}}_{\tau}(f)-1})).$
@@
      If $r\in{\mathfrak{R}}_{\tau}$ and $x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(r)-1}\in{{\rm{dom}}(M)}$, then
      $r^{M}(x_{0},\ldots,x_{{\mathfrak{a}}_{\tau}(r)-1})~{}\Longleftrightarrow~{}r^{N}(\pi(x_{0}),\ldots,\pi(x_{{\mathfrak{a}}_{\tau}(r)-1})).$
@@
    A language $\sigma$ is a sublanguage of a language $\tau$ if $\mathfrak{C}_{\sigma}\subseteq\mathfrak{C}_{\tau}$, $\mathfrak{F}_{\sigma}\subseteq\mathfrak{F}_{\tau}$, $\mathfrak{R}_{\sigma}\subseteq\mathfrak{R}_{\tau}$ and ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\tau}\restriction(\mathfrak{F}_{\sigma}\cup\mathfrak{R}_{\sigma})$.
@@
    satisfying $u[\mathfrak{C}_{\sigma}]=\mathfrak{C}_{\tau}$, $u[\mathfrak{F}_{\sigma}]=\mathfrak{F}_{\tau}$, $u[\mathfrak{R}_{\sigma}]=\mathfrak{R}_{\tau}$ and ${\mathfrak{a}}_{\sigma}(s)={\mathfrak{a}}_{\tau}(u(s))$ for all $s\in\mathfrak{F}_{\sigma}\cup\mathfrak{R}_{\sigma}$.
@@
In [5, Section 4], it is verified that Makowsky’s Theorem 1.1 holds true with respect to the above definitions. In another direction, an unpublished result of Stavi shows that the validity of Vopěnka’s Principle is also equivalent to the fact that analogs of the Downward Löwenheim–Skolem Theorem hold for all abstract logics (see [15, Theorem 6]). For sake of completeness and to motivate the notions introduced in Items (8) and (9) of Definition 5.3, we present a proof of Stavi’s result.
@@
  • Given is a language $\tau$, $M\in Str_{\tau}$ and a binary relation symbol $r$ in $\tau$, then $M\models_{\mathcal{L}_{\Gamma}}\phi_{r}$ holds if and only if $\langle{{\rm{dom}}(M)},r^{M}\rangle$ is a graph that is isomorphic to a graph in $\Gamma$.
Our assumption now yields an uncountable cardinal $\kappa$ with the property that for every language $\tau$ with $|{{\mathfrak{S}}_{\tau}}|<\kappa$, every $\tau$-structure $N$ and every $\phi\in\mathcal{L}_{\Gamma}(\tau)$ with $N\models_{\mathcal{L}_{\Gamma}}\phi$, there exists a substructure $M$ of $N$ with $|{{{\rm{dom}}(M)}}|<\kappa$ and $M\models_{\mathcal{L}_{\Gamma}}\phi$. Since $\Gamma$ is a proper class, we can fix a graph $G$ in $\Gamma$ that has cardinality greater than $\kappa$. Let $\sigma$ denote a language with a single binary relation symbol $e$ and let $\tau$ be a language extending $\sigma$ that adds functions symbols for first-order Skolem functions to $\sigma$. View $G$ as a $\sigma$-structure and let $G^{\prime}$ be a $\tau$-structure whose $\sigma$-reduct is $G$ and that interprets the new function symbols as the corresponding first-order Skolem functions. Then Item (4) in Definition 5.2 ensures that $G^{\prime}\models_{\mathcal{L}_{\Gamma}}\phi_{e}$ holds and, since $|{{\mathfrak{S}}_{\tau}}|<\kappa$, we can use our assumption to find a substructure $H^{\prime}$ of $G^{\prime}$ with $|{H^{\prime}}|<\kappa$ and $H^{\prime}\models_{\mathcal{L}_{\Gamma}}\phi_{e}$. Let $H$ denote the $\sigma$-reduct of $H^{\prime}$. Our setup now ensures that $H$ is a first-order elementary submodel of $G$ and $H\models_{\mathcal{L}_{\Gamma}}\phi_{e}$ holds. This shows that $\langle{{\rm{dom}}(H)},e^{H}\rangle$ is a graph that is isomorphic to an element $K$ of $\Gamma$. Since ${{\rm{dom}}(H)}$ has cardinality less than $\kappa$, we know that $K\neq G$ and $K$ embeds into $G$. These computations show that (1) holds.
@@ 6. Structural properties of abstract logics
(1) Note that, given a language $\sigma$, an application of (6) in Definition 5.2 to the trivial renaming of $\sigma$ into itself ensures that the identity function on $\mathcal{L}(\sigma)$ is the unique permutation $\iota$ of $\mathcal{L}(\sigma)$ with the property that
$M\models_{\mathcal{L}}\phi~{}\Longleftrightarrow~{}M\models_{\mathcal{L}}\iota(\phi)$
@@
(2) Let $\mu$ be a cardinal as in (1), let $\sigma$ be a language in ${\rm{H}}(o_{\mathcal{L}})$, let $u$ be a renaming of $\sigma$ into a language $\tau$ and let $\varpi$ be a bijection between $\mathcal{L}(\sigma)$ and $\mathcal{L}(\tau)$ such that (C) holds. Assume, towards a contradiction that $\varpi\neq u_{*}$. Then $\pi=(u_{*})^{{-}1}\circ\varpi$ is a non-trivial permutation of $\mathcal{L}(\sigma)$ and, since $M_{\sigma,\pi}$ is an element of ${\rm{H}}(\mu)$, our assumptions on $\varpi$ ensure that
@@
holds for all $M\in Str_{\tau_{1}}\cap X$ and all $\phi\in\mathcal{L}(\tau_{1})$. Since ${\rm{H}}(\mu)\subseteq X$, we can now apply (2) to conclude that $\varpi=u_{*}$. ∎
@@
(4) Let $\mu$ be a cardinal as in (1), let $\kappa>\mu$ be a $C^{(n)}$-weakly shrewd cardinal and let $U$ be a ${<}\kappa$-satisfiable $\mathcal{L}$-theory of cardinality $\kappa$. Since $o_{\mathcal{L}}<\kappa$, we can find a language $\sigma$ with $|{{\mathfrak{S}}_{\sigma}}|\leq\kappa$ and $U\subseteq\mathcal{L}(\sigma)$. Then there is a renaming $u$ of $\sigma$ into a language $\tau$ with ${\mathfrak{S}}_{\tau}\subseteq{\rm{H}}(\kappa)$. Set $T=u_{*}[U]$. Then (6) in Definition 5.2 ensures that $T$ is also ${<}\kappa$-satisfiable. Now, pick a cardinal $\kappa<\theta\in C^{(n+2)}$ with $T\in{\rm{H}}(\theta)$ and use (3) to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\mu<\bar{\kappa}<\min(\kappa,\bar{\theta})$, a sublanguage $\bar{\tau}$ of $\tau$ with $|{{\mathfrak{S}}_{\bar{\tau}}}|\leq\bar{\kappa}$ and ${\mathfrak{S}}_{\bar{\tau}}={\mathfrak{S}}_{\tau}\cap{\rm{H}}(\bar{\kappa})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}\cup\{\bar{\kappa},\bar{\tau},\mathcal{L}(\bar{\tau}),z\}\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$, $j(z)=z$, $j(\bar{\tau})=\tau$, $j(\mathcal{L}(\bar{\tau}))=\mathcal{L}(\tau)$, $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$ and $T\in{{\rm{ran}}(j)}$. Pick $\bar{T}\in X$ with $j(\bar{T})=T$. Then $\bar{T}$ is a subset of $\mathcal{L}(\bar{\tau})$ of cardinality $\bar{\kappa}$ and the fact that $\bar{\kappa}\subseteq X$ ensures that $\bar{T}\subseteq X$. Moreover, since $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$, it follows that $\bar{T}=j[\bar{T}]\subseteq T$ and hence $\bar{T}$ is satisfiable. This shows that there is a $\bar{\tau}$-structure $M$ with the property that $\psi(M,\phi,z)$ holds for all $\phi$ in $\bar{T}$. The correctness properties of ${\rm{H}}(\bar{\theta})$ now ensure that this statement holds in $X$ and hence we know that, in ${\rm{H}}(\theta)$, there is a $\tau$-structure $N$ with the property that $\psi(N,\phi,z)$ holds for all $\phi$ in $T$. Using the correctness properties of ${\rm{H}}(\theta)$, we can now conclude that $T$ is satisfiable and, by (6) in Definition 5.2, we also know that $U$ is satisfiable.
@@
(5) Let $\mu$ be a cardinal as in (1), let $\kappa>\mu$ be a $C^{(n)}$-weakly shrewd cardinal, let $\sigma$ be a language with the property that $|{{\mathfrak{S}}_{\sigma}}|<\kappa$, let $I$ be a $\sigma$-structure with $|{{{\rm{dom}}(I)}}|=\kappa$ and let $\phi$ be an element of $\mathcal{L}(\sigma)$ with $I\models_{\mathcal{L}}\phi$. Let $u$ be a renaming of $\sigma$ into a language $\tau\in{\rm{H}}(\kappa)$. Set $K=u^{*}(I)$ and $\phi_{1}=u_{*}(\phi)$. Then $K\models_{\mathcal{L}}\phi_{1}$ and there is an isomorphism $\pi$ between $K$ and a $\tau$-structure $N$ with ${{\rm{dom}}(N)}=\kappa$ and $N\models_{\mathcal{L}}\phi_{1}$. Fix a cardinal $\kappa<\theta\in C^{(n+2)}$. Then $K$ is an element of ${\rm{H}}(\theta)$ and we can use (3) to find a cardinal $\bar{\theta}\in C^{(n)}$, a cardinal $\mu<\bar{\kappa}<\min(\kappa,\bar{\theta})$, a sublanguage $\bar{\tau}$ of $\tau$ with $|{{\mathfrak{S}}_{\bar{\tau}}}|\leq\bar{\kappa}$ and ${\mathfrak{S}}_{\bar{\tau}}={\mathfrak{S}}_{\tau}\cap{\rm{H}}(\bar{\kappa})$, an elementary submodel $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}\cup\{\bar{\kappa},\bar{\tau},\mathcal{L}(\bar{\tau}),z\}\subseteq X$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$, $j(z)=z$, $j(\bar{\tau})=\tau$, $j(\mathcal{L}(\bar{\tau}))=\mathcal{L}(\tau)$, $j\restriction(\mathcal{L}(\bar{\tau})\cap X)={\rm{id}}_{\mathcal{L}(\bar{\tau})\cap X}$ and $N,\phi_{1}\in{{\rm{ran}}(j)}$. The fact that $\tau$ is an element of ${\rm{H}}(\kappa)$ then implies that $\tau=\bar{\tau}\in{\rm{H}}(\bar{\kappa})$, ${\mathfrak{S}}_{\tau}\subseteq X$ and $j\restriction{\mathfrak{S}}_{\tau}={\rm{id}}_{{\mathfrak{S}}_{\tau}}$. Pick $M\in X$ with $j(M)=N$. Then $M$ is a $\tau$-structure with ${{\rm{dom}}(M)}=\bar{\kappa}$ and it follows that $M$ is a substructure of $N$. Next, pick $\phi_{0}$ in $X$ with $j(\phi_{0})=\phi_{1}$. Then $\phi_{0}\in\mathcal{L}(\tau)$ and the fact that $j\restriction(\mathcal{L}(\tau)\cap X)={\rm{id}}_{\mathcal{L}(\tau)\cap X}$ implies that $\phi_{0}=\phi_{1}\in X$. Moreover, using the elementarity of $j$ and the correctness properties of ${\rm{H}}(\theta)$ and ${\rm{H}}(\bar{\theta})$, we know that $M\models_{\mathcal{L}}\phi_{1}$ holds. Then $\pi^{{-}1}\restriction\bar{\kappa}$ is an isomorphism between $M$ and a substructure $J$ of $K$, and we can apply (5) in Definition 5.2 to show that $J\models_{\mathcal{L}}\phi_{1}$. Finally, we can use (5) in Definition 5.2 to conclude that $H=(u^{*})^{{-}1}(J)$ is a substructure of $I$ with $|{{{\rm{dom}}(H)}}|<\kappa$ and $H\models_{\mathcal{L}}\phi$.
@@
Fix $\phi\in\mathcal{L}(\tau)$. By (7) in Definition 5.2, there is a sublanguage $\tau_{0}$ of $\tau$ with $|{{\mathfrak{S}}_{\tau_{0}}}|<o_{\mathcal{L}}$ and $\phi\in\mathcal{L}(\tau_{0})$. Since $|{{\mathfrak{S}}_{\tau}}|^{{<}o_{\mathcal{L}}}\leq\bar{\kappa}\subseteq X$, it follows that $\tau_{0}$ is an element of ${\rm{H}}(\bar{\kappa})\cap X$ and this implies that $j(\tau_{0})=\tau_{0}$ and $j(\mathcal{L}(\tau_{0}))=\mathcal{L}(\tau_{0})$. Now, we can find a language $\tau_{1}\in{\rm{H}}(o_{\mathcal{L}})$ and a renaming $v$ of $\tau_{1}$ into $\tau_{0}$ that is an element of $X$. Our setup then ensures that $\mathcal{L}(\tau_{1})\in{\rm{H}}(\mu)\subseteq X$ and $j(\mathcal{L}(\tau_{1}))=\mathcal{L}(\tau_{1})$. Since (2) shows that $v_{*}$ is the unique bijection between $\mathcal{L}(\tau_{1})$ and $\mathcal{L}(\tau_{0})$ with the property that (C) holds for all $\tau_{1}$-structures $M$ in ${\rm{H}}(\mu)$ and every $\phi\in\mathcal{L}(\tau_{1})$, the fact that all parameters of this existence statement are elements of $X$ that are fixed by $j$ ensures that $v_{*}$ is an element of $X$. In particular, since $\mathcal{L}(\tau_{1})\subseteq{\rm{H}}(\mu)\subseteq X$, we know that $\mathcal{L}(\tau_{0})\subseteq X$ and hence $\phi$ is an element of $X$. ∎